OpenStudy (anonymous):

A store displays seven computers on a shelf side–by–side. Find the total length of the computers on the table if the first computer was six inches wide and the width of each successive computer is three inches longer than the previous one. Show all your work and describe the sigma notation used in answering the question above. please i need help i did everything but this question because i dont understand it and im taking a test right now

OpenStudy (anonymous):

6+3*0 6+3*1 6+3*2 6+3*3\[\sum _{n=0}^6 (6+3*n)=105 \]

OpenStudy (anonymous):

thank you so much